OpenStudy (anonymous):

a cell phone company orders 500 new phones from a manufacturer. if the probability of a phone being defective is 2.6%, predict how many of the phones are likely to be defective. round your answer to the nearest whole number.

pooja195 (pooja195):

Multiply 500 times 2.6% (which is equal to 2.6 / 100)

OpenStudy (helpsos):

use this formula to get your answer= x/500 * 100= 2.6 \[x/500 * 100 = 2.6\] \[x/500 = 2.6/100\] \[x/500 = 0.026\] \[x= 0.026 * 500\] \[x = 13\]

OpenStudy (helpsos):

just want to double check, what do you guys think?

OpenStudy (helpsos):

annyways, hope this helped :)

OpenStudy (anonymous):

yeah thank you so much :)

OpenStudy (helpsos):

np :D